# Some Results on Finitely Splitting Subtrees of Aronszajn Trees

### v1

@@ 4. Promises and Generalized Promises
Let $T$ be an $\omega_{1}$-tree. Let $m<\omega$ be positive and $\gamma<\omega_{1}$. A generalized promise on $T$ with dimension $m$ and base level $\gamma$ is a function $\mathcal{U}$ with domain $[\gamma,\omega_{1})$ satisfying that, for some derived tree $T_{\vec{x}}$ of $T$, for all $\gamma\leq\beta<\omega_{1}$:
@@
    $\mathcal{U}(\beta)$ is a non-empty countable subset of $[T_{\vec{x}}\cap T_{\beta}^{m}]^{\omega}$;
@@
Note that in (2) above, it suffices that this statement holds for $\mathcal{U}(\gamma)$. For assume this and suppose that $\gamma<\beta<\omega_{1}$ and $\mathcal{Y}\in\mathcal{U}(\beta)$. Then by (3), $\mathcal{Y}\upharpoonright\gamma\in\mathcal{U}(\gamma)$. Let $t\subseteq T_{\beta}$ be finite. Then $t\upharpoonright\gamma\subseteq T_{\gamma}$ is finite, so we can find some $\vec{x}_{0}\in\mathcal{Y}\upharpoonright\gamma$ which is disjoint from $t\upharpoonright\gamma$. Now choose $\vec{x}\in\mathcal{Y}$ such that $\vec{x}\upharpoonright\gamma=\vec{x}_{0}$. Then $\vec{x}$ is disjoint from $t$.
A generalized promise $\mathcal{U}$ is essentially an $\omega_{1}$-tree whose elements are sets $\mathcal{Y}$ such that for some $\gamma\leq\beta<\omega_{1}$, $\mathcal{Y}\in[T_{\beta}^{m}]^{\omega}$ is well-distributed, and if $\mathcal{Y}$ and $\mathcal{Z}$ are members of the tree in $[T_{\beta}^{m}]^{\omega}$ and $[T_{\delta}^{m}]^{\omega}$ respectively, where $\beta<\delta$, then $\mathcal{Y}$ is below $\mathcal{Z}$ in the tree iff $\mathcal{Y}=\mathcal{Z}\upharpoonright\beta$. For all $\alpha<\omega_{1}$, level $\alpha$ of this tree is equal to $\mathcal{U}(\gamma+\alpha)$. The union of any cofinal branch of this tree is the tail of some promise on $T$.
@@
Let $T$ be an $\omega_{1}$-tree. Let $m<\omega$ be positive and $\zeta<\omega_{1}$. Suppose that $U$ is a promise with dimension $m$ and base level $\zeta$. Then for some countable ordinal $\gamma<\omega_{1}$ greater than $\zeta$, the function $\mathcal{U}$ with domain $[\gamma,\omega_{1})$ defined by $\mathcal{U}(\beta)=\{U\cap T_{\beta}^{m}\}$ is a generalized promise with dimension $m$ and base level $\gamma$.
@@ 5. Splitting Subtree Functions
For Sections 5-8 and 10, we fix $T$ and $\mathcal{S}$ satisfying that $T$ is a normal Aronszajn tree, $\mathcal{S}\subseteq\omega\setminus 2$, and for any $x\in T$, $|\mathrm{Imm}_{T}(x)|\geq\sup(\mathcal{S})$. Let $s=\min(\mathcal{S})$. In the next section we introduce a forcing which adds an uncountable downwards closed normal subtree $U$ of $T$ such that for all $x\in U$, $|\mathrm{Imm}_{U}(x)|\in\mathcal{S}$. We are mainly interested in two cases: when $\mathcal{S}=\omega\setminus 2$, so $T$ is infinitely splitting and we add a finitely splitting subtree of $T$, and when $\mathcal{S}=\{s\}$ and we add an $s$-splitting subtree of $T$.
@@
If $F$ is a subtree function with top level $\beta$ and $U$ is a promise with dimension $m$ and base level less than $\beta$, then we say that $F$ fulfills $U$ if for every finite set $t\subseteq T_{\beta}$ there exists some $(x_{0},\ldots,x_{m-1})\in U\cap T_{\beta}^{m}$ disjoint from $t$ such that $F(x_{i})=1$ for all $i<m$. This definition of fulfillment has as a precursor Shelah’s notion of fulfillment for his specializing function of [She98, Chapter V, Definition 6.4], and is similar to that used by Moore and Lamei Ramandi in their subtree forcing of [LRM18, Remark 5.2]. The following definition broadens the notion of fulfillment in a natural way to generalized promises, and is similar to that of Abraham-Shelah [AS93, Definition 4.1].
@@
An $\mathcal{S}$-splitting subtree function with top level $\beta<\omega_{1}$ is a subtree function with top level $\beta$ satisfying that for all $x\in T\upharpoonright\beta$,
@@
Successor case: Assume that the statement is true for some countable ordinal $\gamma\geq\beta$, and we prove that it is true for $\gamma+1$. So let $X\subseteq T_{\gamma+1}$ be finite with $(\leq\!s)$-to-one drop-downs to $\beta$ such that for all $x\in X$, $F(x\upharpoonright\beta)=1$. By Lemma 4.5, $X$ has $(\leq\!s)$-to-one drop-downs to $\gamma$ and $X\upharpoonright\gamma$ has $(\leq\!s)$-to-one drop-downs to $\alpha$. Obviously, for all $y\in X\upharpoonright\gamma$, $F(y\upharpoonright\alpha)=1$. By the inductive hypothesis, we can fix an $\mathcal{S}$-splitting subtree function $G$ with top level $\gamma$ such that $G(x\upharpoonright\gamma)=1$ for all $x\in X$ and $G$ fulfills every generalized promise in $\Gamma$.
@@
We define by induction an increasing sequence $\langle Z_{n}:n<\omega\rangle$ of finite subsets of $T_{\gamma+1}$. Our inductive hypothesis is that for all $n<\omega$ and $y\in T_{\gamma}$, $y$ has at most $s$-many immediate successors in $Z_{n}$.
@@
Case 2: $i=1$. Consider $\mathcal{U}_{k_{0}}$, $\mathcal{Y}^{k_{0}}_{k_{1}}$, and $t_{k_{2}}$. Let $l$ be the dimension of $\mathcal{U}_{k_{0}}$. Applying the fact that $F_{n}$ fulfills $\mathcal{U}_{k_{0}}$, we can find $\vec{a}=(a_{0},\ldots,a_{l-1})\in\mathcal{Y}^{k_{0}}_{k_{1}}\upharpoonright\delta_{n}$ which is disjoint from $(Z_{n}\cup t_{k_{2}})\upharpoonright\delta_{n}$. Fix $\vec{a}^{+}=(a_{0}^{+},\ldots,a_{l-1}^{+})\in\mathcal{Y}^{k_{0}}_{k_{1}}$ above $\vec{a}$. Since $\vec{a}$ is injective, $\vec{a}^{+}$ has unique drop-downs to $\delta_{n}$. Note that $\vec{a}^{+}$ is disjoint from $t_{k_{2}}$. Define $Z_{n+1}=Z_{n}\cup\{a_{0}^{+},\ldots,a_{l-1}^{+}\}$. By the inductive hypothesis and the fact that $\{a_{0}^{+},\ldots,a_{l-1}^{+}\}\upharpoonright\delta_{n}$ is disjoint from $Z_{n}\upharpoonright\delta_{n}$, is easily follows that $Z_{n+1}$ which has $(\leq\!s)$-to-one drop-downs to $\delta_{n+1}$
@@
Suppose that $F$ is an $\mathcal{S}$-splitting subtree function with top level $\beta$ and $\mathcal{U}$ is a generalized promise with dimension $l$ and base level $\delta$, where $\delta>\beta$ is a limit ordinal. We say that $\mathcal{U}$ is suitable for $F$ if:
@@
  • there exists a sequence $\langle\vec{b}_{n}:n<\omega\rangle$ of disjoint tuples in $\mathcal{Y}_{0}$, where we let $X_{n}$ be the elements in $\vec{b}_{n}$ for all $n<\omega$;
@@
Suppose that $F$ is an $\mathcal{S}$-splitting subtree function with top level $\beta$, $U$ is a promise on $T$ with base level greater than or equal to $\beta$ and root $(x_{0},\ldots,x_{m-1})$, and for all $i<m$, $F(x_{i}\upharpoonright\beta)=1$. Then for some countable limit ordinal $\delta$ greater than the base level of $U$, the function $\mathcal{U}$ with domain $[\delta,\omega_{1})$ defined by $\mathcal{U}(\beta)=\{U\cap T_{\beta}^{m}\}$ is a generalized promise which is suitable for $F$.
@@
Suppose that $F$ is an $\mathcal{S}$-splitting subtree function with top level $\beta$, $\Gamma$ is a countable collection of generalized promises whose base levels are less than or equal to $\beta$, and $F$ fulfills every member of $\Gamma$. Assume that $\mathcal{U}$ is a generalized promise with base level $\delta>\beta$ which is suitable for $F$. Then there exists a subtree function $F^{+}$ extending $F$ with top level $\delta$ such that $F^{+}$ fulfills every member of $\Gamma\cup\{\mathcal{U}\}$.
@@
$(Z_{n}\cup t_{k_{2}}\cup X_{0}\cup\cdots\cup X_{n+1})\upharpoonright\delta_{n}.$
@@ 6. The Forcing Poset
Since $T$ is normal, we can fix $\xi<N\cap\omega_{1}$ greater than or equal to $\beta$ such that $X$ has unique drop-downs to $\xi$. By Lemma 4.5, $X\upharpoonright\xi$ has $(\leq\!s)$-to-one drop-downs to $\beta$. Applying Proposition 6.2 (Extension) in $N$, find $F^{\prime}$ in $N$ such that $(F^{\prime},\Gamma)\leq(F,\Gamma)$, $(F^{\prime},\Gamma)$ has top level $\xi$, and for all $x\in X$, $F^{\prime}(x\upharpoonright\xi)=1$. Now to simplify notation, let us just assume that $\beta=\xi$ and $F^{\prime}=F$.
@@
Apply Lemma 4.2 to fix a promise $U\subseteq W$ with the same root as $W$, which by elementarity we may assume is in $N$. The root of $U$ is $\vec{x}$, so $\beta$ is both the base level of $U$ and also the top level of $(F,\Gamma)$. Also, $F(x_{i})=1$ for all $i<m$. Applying Lemma 5.6, fix a countable limit ordinal $\delta>\beta$ such that the function $\mathcal{U}$ with domain $[\delta,\omega_{1})$ defined by $\mathcal{U}(\gamma)=\{U\cap T_{\gamma}^{m}\}$ is a generalized promise which is suitable for $F$. By elementarity, we may assume that $\delta$ and $\mathcal{U}$ are in $N$. Apply Proposition 6.3 (Adding Generalized Promises) in $N$ to find some $G$ in $N$ such that $(G,\Gamma\cup\{\mathcal{U}\})\in\mathbb{P}$, $(G,\Gamma\cup\{\mathcal{U}\})\leq(F,\Gamma)$, and $G$ has top level $\delta$.
@@
This complete the induction. Define a condition $(G,\Sigma)$ as follows. Let $\Sigma=\bigcup_{n}\Gamma_{n}$. Define $G$ which extends $\bigcup_{n}F_{n}$ so that for all $x\in T_{N\cap\omega_{1}}$, $G(x)=1$ if $x\in\bigcup_{n}Z_{n}$, and $G(x)=0$ otherwise. We leave it to the reader to check that this works. ∎
@@
Assuming CH, $\mathbb{P}$ is $\omega_{2}$-c.c.
@@ 7. Not Adding Small-Splitting Subtrees
In Section 11, we prove that for any $2<n<\omega$, it is consistent that every $(\geq\!n)$-splitting normal Aronszajn tree contains an uncountable downwards closed $n$-splitting normal subtree, but there exists an infinitely splitting normal Aronszajn tree which contains no uncountable downwards closed $(<\!n)$-splitting subtree. For this purpose, we need the following theorem.
@@
Let $1<n\leq s$. Suppose that $S$ is an $\omega_{1}$-tree which does not contain an uncountable downwards closed $(<\!n)$-splitting subtree. Then $\mathbb{P}$ forces that $S$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree.
@@
Let $1<n\leq s$. Assume that $S$ is an $\omega_{1}$-tree which does not contain an uncountable downwards closed $(<\!n)$-splitting subtree. Let $(F,\Gamma)\in\mathbb{P}$ have top level $\beta$ and suppose that $(F,\Gamma)$ forces that $\dot{W}$ is a downwards closed $(<\!n)$-splitting subtree of $S$.
@@
Define $B$ as the set of all $b\in S$ satisfying:
  • $c\upharpoonright\beta<_{S}b$;
  • the height $\delta$ of $b$ is in $D$;
  • there exists a tuple $\vec{y}=(y_{0},\ldots,y_{l-1})\in T_{\delta}^{l}$ such that $x_{i}<_{T}y_{i}$ for all $i<l$;
  • whenever $(G,\Sigma)\leq(F,\Gamma)$ has some height $\gamma\leq\delta$ and for all $i<l$, $G(y_{i}\upharpoonright\gamma)=1$, then it is not the case that $(G,\Sigma)\Vdash b\notin\dot{W}$.
@@
    since $\dot{W}$ is forced to be downwards closed, for all $\zeta\in D$ with $\beta<\zeta<N\cap\omega_{1}$, $c\upharpoonright\zeta\in B$ as witnessed by $\vec{a}\upharpoonright\zeta$ (since a counter-example to the fourth bullet item can be found in $N$ by elementarity);
@@
    similarly, if $b\in B$ as witnessed by $\vec{y}$, and $\zeta\in D$ is such that $\beta<\zeta\leq\mathrm{ht}_{T}(b)$, then $b\upharpoonright\delta\in B$ as witnessed by $\vec{y}\upharpoonright\zeta$.
@@
Let $\bar{B}$ be the set of $y\in S$ for which there exists some $b\in B$ with $y<_{S}b$. Then $\bar{B}$ is an uncountable downwards closed subtree of $S$. Since $S$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree, $\bar{B}$ is not $(<\!n)$-splitting. So we can fix some $d\in\bar{B}$ which has at least $n$-many immediate successors in $\bar{B}$. Fix distinct immediate successors $d_{0},\ldots,d_{n-1}$ of $d$ in $\bar{B}$. For each $i<n$, choose some $e_{i}\in B$ with $d_{i}<_{S}e_{i}$. By property (3) in the previous paragraph and dropping down to the lowest height of the elements $e_{0},\ldots,e_{n-1}$, we may assume without loss of generality that all of the $e_{i}$’s have the same height $\delta$. For each $i<n$, pick a tuple $\vec{y}^{i}=(y_{0}^{i},\ldots,y_{l-1}^{i})\in T_{\delta}^{l}$ which witnesses that $e_{i}\in B$.
Now for all $i<n$ and $j<l$, $y_{j}^{i}\upharpoonright\beta=x_{j}$, and in particular, $F(y_{j}^{i}\upharpoonright\beta)=1$. Define $X=\{y_{j}^{i}:i<n,\ j<l\}$. Then every member of $T_{\beta}$ either has nothing above it in $X$, or is equal to $x_{j}$ for some $j<l$ and has exactly $n$ many elements above it in $X$, namely, $y_{j}^{0},\ldots,y_{j}^{n-1}$. Since the elements $x_{0},\ldots,x_{l-1}$ are distinct and $n\leq s$, it follows that $X$ has $(\leq\!s)$-to-one drop-downs to $\beta$. As $\delta\in D$, we can fix some $(G,\Sigma)\leq(F,\Gamma)$ with top level $\delta$ such that for all $i<n$ and $j<l$, $G(y_{j}^{i})=1$ and $(G,\Sigma)$ decides $\dot{W}\cap(S\upharpoonright\delta)$. Then for each $i<n$, since $\vec{y}^{i}$ witnesses that $e_{i}\in B$, it is not the case that $(G,\Sigma)\Vdash e_{i}\notin\dot{W}$. But $(G,\Sigma)$ forces that $\dot{W}$ is $(<\!n)$-splitting and $(G,\Sigma)$ decides $\dot{W}\cap S_{\mathrm{ht}_{S}(d)+1}$. So there must exist some $i<n$ such that $(G,\Sigma)\Vdash d_{i}\notin\dot{W}$. As $\dot{W}$ is forced to be downwards closed and $d_{i}<_{S}e_{i}$, $(G,\Sigma)\Vdash e_{i}\notin\dot{W}$, which is a contradiction. ∎

### v2

@@ 4. Promises and Generalized Promises
Let $T$ be an $\omega_{1}$-tree. Let $m<\omega$ be positive and $\gamma<\omega_{1}$. A generalized promise on $T$ with dimension $m$ and base level $\gamma$ is a function $\mathcal{U}$ with domain $\gamma,\omega_{1})$ satisfying that, for some derived tree $T_{\vec{x}}$ of $T$, for all $\gamma\leq\beta<\omega_{1}$:
@@
    $\mathcal{U}(\beta)$ is a non-empty countable collection of infinite subsets of $T_{\vec{x}}\cap T_{\beta}^{m}$;
@@
It is implicit in the above that $\vec{x}$ has size $m$ and its elements have height at most $\gamma$. Note that in (2), it suffices that this statement holds for $\mathcal{U}(\gamma)$. For assume this and suppose that $\gamma<\beta<\omega_{1}$ and $\mathcal{Y}\in\mathcal{U}(\beta)$. Then by (3), $\mathcal{Y}\upharpoonright\gamma\in\mathcal{U}(\gamma)$. Let $t\subseteq T_{\beta}$ be finite. Then $t\upharpoonright\gamma\subseteq T_{\gamma}$ is finite, so we can find some $\vec{b}_{0}\in\mathcal{Y}\upharpoonright\gamma$ which is disjoint from $t\upharpoonright\gamma$. Now choose $\vec{b}\in\mathcal{Y}$ such that $\vec{b}\upharpoonright\gamma=\vec{b}_{0}$. Then $\vec{b}$ is disjoint from $t$.
A generalized promise $\mathcal{U}$ is essentially an $\omega_{1}$-tree whose elements are sets $\mathcal{Y}$ such that for some $\gamma\leq\beta<\omega_{1}$, $\mathcal{Y}\in[T_{\beta}^{m}]^{\omega}$ is well-distributed, and if $\mathcal{Y}$ and $\mathcal{Z}$ are members of the tree in $T_{\beta}^{m}]^{\omega}$ and $T_{\delta}^{m}]^{\omega}$ respectively, where $\beta<\delta$, then $\mathcal{Y}$ is below $\mathcal{Z}$ in the tree iff $\mathcal{Y}=\mathcal{Z}\upharpoonright\beta$. For all $\alpha<\omega_{1}$, level $\alpha$ of this tree is equal to $\mathcal{U}(\gamma+\alpha)$. The union of any cofinal branch of this tree is the tail of some promise on $T$.
@@
Let $T$ be an $\omega_{1}$-tree. Let $m<\omega$ be positive and $\zeta<\omega_{1}$. Suppose that $U$ is a promise with dimension $m$ and base level $\zeta$. Then for some countable ordinal $\gamma<\omega_{1}$ greater than $\zeta$, the function $\mathcal{U}$ with domain $\gamma,\omega_{1})$ defined by $\mathcal{U}(\beta)=\{U\cap T_{\beta}^{m}\}$ is a generalized promise with dimension $m$ and base level $\gamma$.
@@ 5. Splitting Subtree Functions
For Sections 5-8 and 10, we fix $T$ and $\mathcal{S}$ satisfying that $T$ is a normal Aronszajn tree, $\mathcal{S}\subseteq\omega\setminus 2$ is non-empty, and for any $x\in T$, $|\mathrm{Imm}_{T}(x)|\geq\sup(\mathcal{S})$. Let $s=\min(\mathcal{S})$. In the next section we introduce a forcing which adds an uncountable downwards closed normal subtree $U$ of $T$ such that for all $x\in U$, $|\mathrm{Imm}_{U}(x)|\in\mathcal{S}$. We are mainly interested in two cases: when $\mathcal{S}=\omega\setminus 2$, so $T$ is infinitely splitting and we add a finitely splitting subtree of $T$, and when $\mathcal{S}=\{s\}$ and we add an $s$-splitting subtree of $T$.
@@
If $F$ is a subtree function with top level $\beta$ and $U$ is a promise with dimension $m$ and base level less than $\beta$, then we say that $F$ fulfills $U$ if for every finite set $t\subseteq T_{\beta}$ there exists some $(x_{0},\ldots,x_{m-1})\in U\cap T_{\beta}^{m}$ disjoint from $t$ such that $F(x_{i})=1$ for all $i<m$. This definition of fulfillment has as a precursor Shelah’s notion of fulfillment for his specializing function of [She98, Chapter V, Definition 6.4], and is similar to that used by Lamei Ramandi and Moore in their subtree forcing of [LRM18, Remark 5.2]. The following definition broadens the notion of fulfillment in a natural way to generalized promises, and is similar to that of Abraham-Shelah [AS93, Definition 4.1].
@@
An $\mathcal{S}$-splitting subtree function with top level $\beta<\omega_{1}$ is any subtree function $F$ with top level $\beta$ satisfying that for all $x\in T\upharpoonright\beta$, if $F(x)=1$, then
@@
Successor case: Assume that the statement is true for some countable ordinal $\gamma\geq\beta$, and we prove that it is true for $\gamma+1$. So let $X\subseteq T_{\gamma+1}$ be finite with $(\leq\!s)$-to-one drop-downs to $\beta$ such that for all $x\in X$, $F(x\upharpoonright\beta)=1$. By Lemma 4.5, $X$ has $(\leq\!s)$-to-one drop-downs to $\gamma$ and $X\upharpoonright\gamma$ has $(\leq\!s)$-to-one drop-downs to $\beta$. Obviously, for all $y\in X\upharpoonright\gamma$, $F(y\upharpoonright\beta)=1$. By the inductive hypothesis, we can fix an $\mathcal{S}$-splitting subtree function $G$ with top level $\gamma$ such that $G(x\upharpoonright\gamma)=1$ for all $x\in X$ and $G$ fulfills every generalized promise in $\Gamma$.
@@
We define by induction an increasing sequence $\langle Z_{n}:n<\omega\rangle$ of finite subsets of $T_{\gamma+1}$. Our inductive hypothesis is that for all $n<\omega$ and $\in T_{\gamma}, yy$ has at most $s$-many immediate successors in $Z_{n}$.
@@
Case 2: $i=1$. Consider $\mathcal{U}_{k_{0}}$, $\mathcal{Y}^{k_{0}}_{k_{1}}$, and $t_{k_{2}}$. Let $l$ be the dimension of $\mathcal{U}_{k_{0}}$. Applying the fact that $F_{n}$ fulfills $\mathcal{U}_{k_{0}}$, we can find $\vec{a}=(a_{0},\ldots,a_{l-1})\in\mathcal{Y}^{k_{0}}_{k_{1}}\upharpoonright\delta_{n}$ which is disjoint from $(Z_{n}\cup t_{k_{2}})\upharpoonright\delta_{n}$. Fix $\vec{a}^{+}=(a_{0}^{+},\ldots,a_{l-1}^{+})\in\mathcal{Y}^{k_{0}}_{k_{1}}$ above $\vec{a}$. Since $\vec{a}$ is injective, $\vec{a}^{+}$ has unique drop-downs to $\delta_{n}$. Note that $\vec{a}^{+}$ is disjoint from $t_{k_{2}}$. Define $Z_{n+1}=Z_{n}\cup\{a_{0}^{+},\ldots,a_{l-1}^{+}\}$. By the inductive hypothesis and the fact that $\{a_{0}^{+},\ldots,a_{l-1}^{+}\}\upharpoonright\delta_{n}$ is disjoint from $Z_{n}\upharpoonright\delta_{n}$, it easily follows that $Z_{n+1}$ has $(\leq\!s)$-to-one drop-downs to $\delta_{n+1}$.
@@
Suppose that $F$ is an $\mathcal{S}$-splitting subtree function with top level $\beta$ and $\mathcal{U}$ is a generalized promise with dimension $l$ and base level $\delta$, where $\delta>\beta$ is a limit ordinal. We say that $\mathcal{U}$ is suitable for $F$ if for some sequences $\langle\delta_{n}:n<\omega\rangle$ and $\langle\vec{b}_{n}:n<\omega\rangle$:
@@
  • $\langle\delta_{n}:n<\omega\rangle$ is an increasing sequence of countable ordinals with supremum $\delta$, where $\delta_{0}>\beta$;
  • $\langle\vec{b}_{n}:n<\omega\rangle$ is a sequence of disjoint tuples in $\mathcal{Y}_{0}$, where we let $X_{n}$ be the elements in $\vec{b}_{n}$ for all $n<\omega$;
@@
Suppose that $F$ is an $\mathcal{S}$-splitting subtree function with top level $\beta$, $U$ is a promise on $T$ with base level greater than or equal to $\beta$ and root $(x_{0},\ldots,x_{m-1})$, and for all $i<m$, $F(x_{i}\upharpoonright\beta)=1$. Then for some countable limit ordinal $\delta$ greater than the base level of $U$, the function $\mathcal{U}$ with domain $\delta,\omega_{1})$ defined by $\mathcal{U}(\beta)=\{U\cap T_{\beta}^{m}\}$ is a generalized promise which is suitable for $F$.
@@
Suppose that $F$ is an $\mathcal{S}$-splitting subtree function with top level $\beta$, $\Gamma$ is a countable collection of generalized promises whose base levels are less than or equal to $\beta$, and $F$ fulfills every member of $\Gamma$. Assume that $\mathcal{U}$ is a generalized promise with base level $\delta>\beta$ which is suitable for $F$. Then there exists an $\mathcal{S}$-splitting subtree function $F^{+}$ extending $F$ with top level $\delta$ such that $F^{+}$ fulfills every member of $\Gamma\cup\{\mathcal{U}\}$.
@@
$(Z_{n}\cup t_{k_{2}}\cup X_{0}\cup\cdots\cup X_{n+1})\upharpoonright\delta_{n}$
@@ 6. The Forcing Poset
Since $T$ is normal, we can fix $\xi<N\cap\omega_{1}$ greater than or equal to $\beta$ such that $X$ has unique drop-downs to $\xi$. By Lemma 4.5, $X\upharpoonright\xi$ has $(\leq\!s)$-to-one drop-downs to $\beta$. Applying Proposition 6.2 (Extension) in $N$, find $F^{\prime}$ in $N$ such that $(F^{\prime},\Gamma)\leq(F,\Gamma)$, $(F^{\prime},\Gamma)$ has top level $\xi$, and for all $x\in X$, $F^{\prime}(x\upharpoonright\xi)=1$. Now to simplify notation, let us just assume that $\beta=\xi$ and $F=F^{\prime}$.
@@
Apply Lemma 4.2 to fix a promise $U\subseteq W$ with the same root as $W$, which by elementarity we may assume is in $N$. The root of $U$ is $\vec{x}$, so $\beta$ is both the base level of $U$ and also the top level of $(F,\Gamma)$. Also, $F(x_{i})=1$ for all $i<m$. Applying Lemma 5.6, fix a countable limit ordinal $\delta>\beta$ such that the function $\mathcal{U}$ with domain $\delta,\omega_{1})$ defined by $\mathcal{U}(\gamma)=\{U\cap T_{\gamma}^{m}\}$ is a generalized promise which is suitable for $F$. By elementarity, we may assume that $\delta$ and $\mathcal{U}$ are in $N$. Apply Proposition 6.3 (Adding Generalized Promises) in $N$ to find some $G$ in $N$ such that $(G,\Gamma\cup\{\mathcal{U}\})\in\mathbb{P}$, $(G,\Gamma\cup\{\mathcal{U}\})\leq(F,\Gamma)$, and $G$ has top level $\delta$.
@@
This completes the induction. Define a condition $(G,\Sigma)$ as follows. Let $\Sigma=\bigcup_{n}\Gamma_{n}$. Define $G$ which extends $\bigcup_{n}F_{n}$ so that for all $x\in T_{N\cap\omega_{1}}$, $G(x)=1$ if $x\in\bigcup_{n}Z_{n}$, and $G(x)=0$ otherwise. We leave it to the reader to check that this works. ∎
@@
It is routine to show that whenever $G$ is a generic filter on $\mathbb{P}$, then in $V[G]$ the set
$U=\{x\in T:\exists(F,\Gamma)\in G\ F(x)=1\}$
is an uncountable downwards closed subtree of $T$ which is normal and satisfies that for all $x\in U$, $|\mathrm{Imm}_{U}(x)|\in\mathcal{S}$.
@@
Assuming CH, $\mathbb{P}$ is $\omega_{2}$-$c.$
@@ 7. Not Adding Small-Splitting Subtrees
Recall that $T$ is a fixed normal Aronszajn tree, $\mathcal{S}\subseteq\omega\setminus 2$, for all $x\in T$, $|\mathrm{Imm}_{T}(x)|\geq\sup(\mathcal{S})$, and $\mathbb{P}=\mathbb{P}(T,\mathcal{S})$. In Section 11 we prove that for any $2<n<\omega$, it is consistent that every $(\geq\!n)$-splitting normal Aronszajn tree contains an uncountable downwards closed $n$-splitting normal subtree, but there exists an infinitely splitting normal Aronszajn tree which contains no uncountable downwards closed (<n)(<\!n)-splitting subtree. For this purpose, we need the following theorem.
@@
Let $1<n\leq s$. Suppose that $S$ is an $\omega_{1}$-tree which does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree. Then $\mathbb{P}$ forces that $S$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree.
@@
Let $1<n\leq s$. Assume that $S$ is an $\omega_{1}$-tree which does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree. Let $(F,\Gamma)\in\mathbb{P}$ have top level $\beta$ and suppose that $(F,\Gamma)$ forces that $\dot{W}$ is a downwards closed (<n)(<\!n)-splitting subtree of $S$.
@@
Define $B$ as the set of all $b\in S$ satisfying that $c\upharpoonright\beta<_{S}b$, the height $\delta$ of $b$ is in $D$, and there exists a tuple $\vec{y}=(y_{0},\ldots,y_{l-1})\in T_{\delta}^{l}$ such that $x_{i}<_{T}y_{i}$ for all $i<l$, and whenever $(G,\Sigma)\leq(F,\Gamma)$ has some top level $\gamma\leq\delta$ and for all $i<l$, $G(y_{i}\upharpoonright\gamma)=1$, then it is not the case that $(G,\Sigma)\Vdash b\notin\dot{W}$.
@@
    since $\dot{W}$ is forced to be downwards closed, for all $\zeta\in D$ with $\beta<\zeta<N\cap\omega_{1}$, $c\upharpoonright\zeta\in B$ as witnessed by $\vec{a}\upharpoonright\zeta$;
@@
    similarly, if $b\in B$ as witnessed by $\vec{y}$, and $\zeta\in D$ is such that $\beta<\zeta\leq\mathrm{ht}_{T}(b)$, then $b\upharpoonright\zeta\in B$ as witnessed by $\vec{y}\upharpoonright\zeta$.
@@
Let $\bar{B}$ be the set of $y\in S$ for which there exists some $b\in B$ with $y<_{S}b$. Then $\bar{B}$ is an uncountable downwards closed subtree of $S$. Since $S$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree, $\bar{B}$ is not (<n)(<\!n)-splitting. So we can fix some $d\in\bar{B}$ which has at least $n$-many immediate successors in $\bar{B}$. Fix distinct immediate successors $d_{0},\ldots,d_{n-1}$ of $d$ in $\bar{B}$. For each $i<n$, choose some $e_{i}\in B$ with $d_{i}<_{S}e_{i}$. By property (3) in the previous paragraph and dropping down to the lowest height of the elements $e_{0},\ldots,e_{n-1}$, we may assume without loss of generality that all of the $e_{i}$’s have the same height $\delta$. For each $i<n$, pick a tuple $\vec{y}^{i}=(y_{0}^{i},\ldots,y_{l-1}^{i})\in T_{\delta}^{l}$ which witnesses that $e_{i}\in B$.
Now for all $i<n$ and $j<l$, $y_{j}^{i}\upharpoonright\beta=x_{j}$, and in particular, $F(y_{j}^{i}\upharpoonright\beta)=1$. Define $X=\{y_{j}^{i}:i<n,\ j<l\}$. Then every member of $T_{\beta}$ either has nothing above it in $X$, or is equal to $x_{j}$ for some $j<l$ and has exactly the elements $y_{j}^{0},\ldots,y_{j}^{n-1}$ above it in $X$. Since $n\leq s$, it follows that $X$ has $(\leq\!s)$-to-one drop-downs to $\beta$. As $\delta\in D$, we can fix some $(G,\Sigma)\leq(F,\Gamma)$ with top level $\delta$ such that for all $i<n$ and $j<l$, $G(y_{j}^{i})=1$ and $(G,\Sigma)$ decides $\dot{W}\cap(S\upharpoonright\delta)$. Then for each $i<n$, since $\vec{y}^{i}$ witnesses that $e_{i}\in B$, it is not the case that $(G,\Sigma)\Vdash e_{i}\notin\dot{W}$. But $(G,\Sigma)$ forces that $\dot{W}$ is (<n)(<\!n)-splitting and $(G,\Sigma)$ decides $\dot{W}\cap S_{\mathrm{ht}_{S}(d)+1}$. So there must exist some $i<n$ such that $(G,\Sigma)\Vdash d_{i}\notin\dot{W}$. As $\dot{W}$ is forced to be downwards closed and $d_{i}<_{S}e_{i}$, $(G,\Sigma)\Vdash e_{i}\notin\dot{W}$, which is a contradiction. ∎